# GenEscape: Hierarchical Multi-Agent Generation of Escape Room Puzzles

### v1

@@ Abstract
GenEscape: Hierarchical Multi-Agent Generation of Escape Room Puzzles
Mengyi Shan  Brian Curless  Ira Kemelmacher-Shlizerman  Steve Seitz University of Washington Seattle, WA {shanmy,curless,kemelmi,seitz}@cs.washington.edu
@@ 1 Introduction
Escape rooms are environments designed as interactive puzzles, where players must explore a confined scene, manipulate objects in a precise order, and ultimately exit the room. We challenge modern Vision-Language Models (VLMs) with the task of designing and building 2D escape room image puzzles. While those models produce aesthetically compelling images OpenAI (2024); DeepMind (2024); Wang et al. (2024a); Li et al. (2024), they struggle with complex scenes that require fine-grained spatial relationships, physical affordance reasoning, or multi-step functional coherence.
@@
We propose a novel multi-agent  Zhou et al. (2024); Wang et al. (2024b); Li et al. (2025) interaction framework for escape room image generation. As in Fig 1, each agent contributes to a specific aspect of the scene (puzzle design, object placement, spatial consistency, or visual affordance verification) through iterative communication and refinement. This division of labor allows the system to reason about puzzle structure and object semantics in a modular way without sacrificing visual quality.
@@ 2 Related Work
Prompt Optimization. Prompt optimization is a powerful strategy for guiding LLM without weight updates. In NLP, techniques such as Shin et al. (2020); Lester et al. (2021); Ouyang et al. (2022) show that well-crafted prompts improve downstream performance. Recent studies explore black-box prompt optimization Diao et al. (2023); Zhou and Neubig (2023); Park et al. (2023) for tasks like QA and reasoning. In vision, prompt engineering and multi-turn refinement help steer diffusion and vision-language models Liu et al. (2022); Hertz et al. (2023); Hao et al. (2023); Mo et al. (2024); Wu et al. (2024). Our approach introduces a hierarchical framework with symbolic intermediates, reducing local minima risks and improving convergence.
Multi-Agent Systems. Recent work has explored how multiple specialized agents can collaboratively solve complex tasks in language and vision Zhang et al. (2024b); Xiong et al. (2025); Google Research (2024). In visual generation, multi-agent setups have been applied to interactive storytelling, scene composition, and instruction-following environments, where agents assume distinct roles such as planning, verification, and rendering Xu et al. (2024); Zhao et al. (2024); Hu et al. (2021); Gao et al. (2024). Most prior work either focuses on interactive dialogue or relies on fixed procedural pipelines.
Generating Puzzles. LLMs have been applied to puzzle generation and solving across multimodal domains. Raihan et al. (2024) combined chain-of-thought prompting with LLMs for word and sentence puzzles. Chia et al. (2024) introduced a diagnostic benchmark of abstract visual puzzles to assess multimodal reasoning. Estermann et al. (2024) proposed a reinforcement learning benchmark targeting logic-based algorithmic reasoning. EscapeCraft Wang et al. (2025) also focuses on escape room generation, but relies on 3D assets, differing from our 2D image-based approach.
@@ 3.3 Scene Graph Optimization
In this stage, we refine the scene graph purely through symbolic reasoning, without involving any visual modality. The Player agent attempts to solve the puzzle by generating a sequence of actions based solely on the scene graph structure. This proposed solution is then evaluated by the Examiner agent against the intended one. The Examiner summarizes key differences in bullet points—for example, highlighting when a player exploits an unintended shortcut. In such cases, the Examiner revises the scene graph to block the shortcut and reinforce the expected solution path. Through iterative correction, the scene graph is adjusted until it supports a coherent and solvable strategy that aligns with the official solution. This process ensures that downstream image generation will faithfully reflect a functionally valid puzzle.
@@
Output : Final image $I$ that supports $S$
@@ 3.4 Affordance-Guided Local Editing
After the scene graph is finalized, we prompt the Builder to generate a 2D layout (Fig.  2 right) where each object is represented by an abstract icon positioned according to the spatial relationships in the scene graph. The same Player–Examiner interaction loop is reused to guarantee that the layout visually supports the intended puzzle solution.
Once the layout is verified, the Builder renders a photorealistic image based on the layout and the description. In this final stage, the Examiner again compares the Player’s actions—now based on the image—to the intended solution. If the Player misinterprets object affordances, the Examiner identifies which visual cues are lacking. Using the icon position from the layout stage, we then apply local image editing to enhance or suppress affordances to steer perception toward the correct interaction.
@@ 4 Experiment
We use GPT-4o  OpenAI (2024) API as our base model and generate square images at 1024 $\times$1024 resolution with the highest quality. We summarize the prompts for each step in the Appendix.
@@ 4.1 Evaluation Metrics
We evaluate our method using three human evaluation metrics. Solvability checks whether a player can infer the intended solution from the image. Shortcut Avoidance measures whether trivial, unintended solutions are blocked. Spatial Alignment assesses how well the visual output reflects the designed object relations. We also report the LongCLIP Zhang et al. (2024a) score measuring semantic similarity between the Designer prompt and the image, and the average number of image API calls (not applicable for vanilla baselines without refinement on images). Evaluation is conducted on 15 scene settings with 2 core objects each. Ten annotators are asked to pick the best image per case for the three evaluation aspects, given the setting and object names; we report the percentage of the results getting picked as best.
@@ 4.2 Analysis
We compare against vanilla GPT-4o OpenAI (2024) and several ablations by removing refinement stages on the scene graph, layout sketch, and final image. Qualitative results are shown in Fig 3. Take the top-right classroom scene with the ladder as an example: naively prompting GPT for an “escape room” (top left) yields a scene with generic visual cues but no clear solution path. Using only the description includes key elements but lacks spatial grounding (e.g., a lock on the blackboard). Adding scene graphs and layouts improves spatial coherence and supports solvability. Final image editing removes shortcuts (e.g., climbable desks) and artifacts (e.g., malformed hooks).
Quantitative results in Tab 1 confirm that the scene graph improves the model’s ability to generate complex spatial relationships involving more than four objects, thus improving solvability; the layout sketch reduces the number of generations; and visual affordance refinement enhances local object characteristics, preventing shortcut solutions and promoting solution alignment. We didn’t see a clear improvement in the CLIP score, likely because the direct editing on images might change objects’ statuses to increase affordance awareness, disobeying the first stage’s prompt.
@@ 5 Conclusion
We presented a hierarchical multi-agent framework for generating escape room puzzles that are both visually realistic and logically solvable. By decomposing the task into symbolic reasoning, layout planning, and affordance-aware image refinement, our method ensures that each generated scene supports a coherent solution path and clear visual cues.
Limitations
@@
Ethical Consideration
While our system is designed for virtual game development, it builds upon large language models (LLMs) that are known for their open-ended creativity and ability to generate realistic, interactive scenarios. When prompted without safeguards, these models may suggest escape room ideas involving hazardous mechanisms or unsafe real-world interactions, such as the use of fire, heavy objects, or confined spaces, that could be misinterpreted or physically replicated. This raises a broader concern: even though the generated content is intended for virtual or fictional use, it may inspire real-world setups that pose risks to safety, especially among unsupervised users or those lacking experience in safe design. Developers and researchers leveraging such generative tools must remain vigilant about unintended consequences, embedding appropriate content filters and clearly communicating that generated outputs are not suitable for real-life execution without expert oversight.
Appendix A Solution to Figure 3
A.1 Pool + Lifebuoy
@@ A.3 Prison + Ice
A.4 Birthday Party + Balloon
@@ A.6 Hospital + Magnifier
Figure 4: Additional Results without usage of locks. Left columns are naive generation (top) and assisted by a textual design (bottom). The second and third columns show the scene graph and 2D layout (top), and results with them (bottom). Right columns are the final results after visual signal-based image optimization.
Appendix B Additional Results
We additionally attach some creative results our model generates. Specifically, we ask the model to generate escape rooms where the human player can exit the room not by unlocking the door. See Fig. 4 for two such samples. Solutions are provided below.
B.1 Amusement Park + Seesaw
@@
    Observe the sign "BALANCE THE FUN TO ESCAPE!" to understand that the seesaw must be balanced to open the exit.
@@
    Examine the right end of the seesaw and identify the 50kg and 20kg weights locked inside a cage.
@@
    Realize that the cage is fixed and those weights cannot be moved—this end of the seesaw is already weighted.
@@
    Collect the 100kg kettlebell and the golden elephant head from the ground near the left side of the seesaw.
@@
    Place the 100kg kettlebell and elephant head carefully on the left side of the seesaw to counterbalance the fixed weights on the right.
@@
    As the seesaw balances, the pressure on the right side lifts, unlocking the hidden trapdoor marked “EXIT”.
  7. 7.
    Pull the handle on the EXIT mat to open the trapdoor and escape the room.
B.2 Basement + Bomb
@@
    Take the crowbar from the wall-mounted shelf.
@@
    Use the crowbar to break the glass dome encasing the bomb.
@@
    Retrieve the matchstick from the open cabinet.
@@
    Light the fuse on the bomb using the matchstick.
@@
    Wait for the bomb to explode and destroy the weakened wall.
@@
    Exit the room through the opening in the destroyed wall.
Appendix C Human Interface
Figure 5: Part 1: Real case of human-agent interaction. The human is typing in possible solutions trying to solve the escape room puzzle, while the AI is providing feedback and gradually guiding the player towards the ideal solution. Followed by the next figure.
Figure 6: Part 2: Real case of human-agent interaction. The human is typing in possible solutions trying to solve the escape room puzzle, while the AI is providing feedback and gradually guiding the player towards the ideal solution.
We build a human interface that lets a human player look at the generated escape room puzzle and try to solve it. Our AI agent, with the official solution in mind, is able to judge whether the human solution works or not, and step by step guides the human player toward the correct solution. See Fig. 5 and Fig. 6 for an example of the game-playing procedure.
Appendix D Sample Prompts
We provide the core, minimal set of prompts used for the multi-agent system. In practice, users can tweak the prompt to achieve certain effects (i.e., generating a certain style, stricter or looser on criticizing, etc.)
D.1 Designer
  • Help me design a mechanical and tactile escape room puzzle. There should be a visible lock on an exit door. And the player should take advantage of different interactive objects in the room to unlock that door. All the objects that are required to solve the puzzle should be visible. (There shouldn’t be things like a key locked in a wooden box that the user cannot spot unless they open the box. ) Answer in a paragraph of scene description and a list of steps to solve it. Omit steps like "looking around" or "noticing xxx". List solid actions. The scene sets in a <scene> with <objects>. The scene is in a mystery but bright and vibrant atmosphere. Try to generate a very easy puzzle that can be solved in 3 steps with at most 4 objects involved.
  • Can you generate a scene graph for this scene in yaml format? It should be represented in a tree form, with the room as the highest ancestor. It should include all objects in the room. If something is attached or stored inside another object, they should be represented as parent and child.
  • Can you generate a minimalist, black and white 2D scene layout sketch for this scene? Make sure the generation is from side view. Use 2D icons to represent central and salient objects in the scene, as well as their correct spatial relationship. Use text as marking if objects have specific conditions.
D.2 Player
If you are a player, how would you utilize existing objects in this scene to exit from this room? Omit steps like observing or surveying and only focus on solid interactions with the objects. Pick one most reasonable sequence of actions.
D.3 Examiner
  • Here are the steps to solve the puzzle: <solution>. Do you think your solution is the same as it? It’s acceptable to have a small difference, but the major logic and order should be the same. If yes, just say yes. If no, summarize the major differences in bullet points between yours and the official solutions.
  • (If answered no to the previous question) What would be some changes to the scene graph that could provide more visual cues towards the correct solution? Rewrite that scene graph by adding or deleting objects, rearranging their spatial relationship, or modifying conditions, with emphasis on crucial objects.
  • What are some modifications to the layout that can make the visual cues motivating the players more for the official solution?
  • How would you edit the image that can make the visual cues motivating the players more for the official solution? Focus on editing on object at a time. Describe the local object you’d like to change, its position, and its expected new condition.
D.4 Builder
  • We are designing an escape room image puzzle. You are supposed to summarize the major objects in the scene and generate an image of 2D layout of the escape room. Draw crucial objects, and write their conditions (if any) next to the objects.
  • We are designing an escape room image puzzle. You are supposed to summarize the major objects in the scene and generate a photorealistic image of the escape room, adhering to <description> and <layout>.
Appendix E Human Evaluation Instructions
You will be shown six generated escape room images based on the same underlying scene graph and puzzle description—each produced by a different model. Your task is to evaluate these results along three criteria and pick the best one for each criterion.
Please consider the following aspects carefully:
  • Solvability: Does the image provide enough visual cues for a player to correctly infer the intended solution? A good image should make the necessary objects, their functions, and the sequence of interactions clear enough that a human could reasonably solve the puzzle.
  • Shortcut Avoidance: Does the image prevent trivial or unintended solutions? For example, if the puzzle is designed to require fixing a stool, the image should not allow the player to simply jump on a desk and bypass the challenge. The best results will clearly block or visually discourage such shortcuts.
  • Graph Alignment: How well does the visual content match the structure and object relationships in the input scene graph? Pay attention to whether key objects are present, in the correct spatial positions, and whether their states (e.g., broken, locked, accessible) are visually reflected in the image.
For each criterion, assign a rank from 1 (best) to 4 (worst). Rankings must be unique (no ties). Please base your judgment on visual reasoning and consistency, not image quality or aesthetics alone.
Appendix F Annotator Details
All human annotators involved in this study are students. They volunteered to participate without compensation and were fully informed that their evaluations would be used as part of a research project to assess the performance of different escape room generation models. Each annotator provided informed consent and was instructed to focus on the functional aspects of the images.

### v2

@@ 1 Introduction
Escape rooms are environments designed as interactive puzzles, where players must explore a confined scene, manipulate objects in a precise order, and ultimately exit the room. We challenge modern Vision-Language Models (VLMs) with the task of designing and building 2D escape room image puzzles. While those models produce aesthetically compelling images [15, 2, 20, 12], they struggle with complex scenes that require fine-grained spatial relationships, physical affordance reasoning, or multi-step functional coherence.
@@
We propose a novel multi-agent  [29, 21, 11] interaction framework for escape room image generation. As in Fig 1, each agent contributes to a specific aspect of the scene (puzzle design, object placement, spatial consistency, or visual affordance verification) through iterative communication and refinement. This division of labor allows the system to reason about puzzle structure and object semantics in a modular way without sacrificing visual quality.
@@ 2 Related Work
Prompt Optimization. Prompt optimization is a powerful strategy for guiding LLM without weight updates. In NLP, techniques such as [19, 10, 16] show that well-crafted prompts improve downstream performance. Recent studies explore black-box prompt optimization [3, 30, 17] for tasks like QA and reasoning. In vision, prompt engineering and multi-turn refinement help steer diffusion and vision-language models [13, 8, 7, 14, 23]. Our approach introduces a hierarchical framework with symbolic intermediates, reducing local minima risks and improving convergence.
Multi-Agent Systems. Recent work has explored how multiple specialized agents can collaboratively solve complex tasks in language and vision [27, 24, 6]. In visual generation, multi-agent setups have been applied to interactive storytelling, scene composition, and instruction-following environments, where agents assume distinct roles such as planning, verification, and rendering [25, 28, 9, 5]. Most prior work either focuses on interactive dialogue or relies on fixed procedural pipelines.
Generating Puzzles. LLMs have been applied to puzzle generation and solving across multimodal domains. Raihan et al. [18] combined chain-of-thought prompting with LLMs for word and sentence puzzles. Chia et al. [1] introduced a diagnostic benchmark of abstract visual puzzles to assess multimodal reasoning. Estermann et al. [4] proposed a reinforcement learning benchmark targeting logic-based algorithmic reasoning. EscapeCraft [22] also focuses on escape room generation, but relies on 3D assets, differing from our 2D image-based approach.
@@ 3.3 Scene Graph Optimization
In this stage, we refine the scene graph purely through symbolic reasoning, without involving any visual modalities. The Player agent attempts to solve the puzzle by generating a sequence of actions based solely on the scene graph structure. This proposed solution is then evaluated by the Examiner agent, who compares it against the intended solution. The Examiner highlights key discrepancies—such as when the player exploits an unintended shortcut—and summarizes them in bullet-point feedback. In response, the Examiner revises the scene graph to eliminate such shortcuts and reinforce the intended solution path. Through iterative refinement, the scene graph is progressively updated until it supports a coherent, logically sound strategy that aligns with the ground-truth solution. This symbolic verification step ensures that downstream image generation faithfully captures a functionally valid and solvable puzzle.
@@
Output : Final image II that supports $S$
@@ 3.4 Affordance-Guided Local Editing
After the scene graph is finalized, we prompt the Builder to generate a 2D layout (Fig.  2 right) where each object is represented by an abstract icon positioned according to the spatial relationships in the scene graph. The same Player–Examiner interaction loop is reused to guarantee that the layout visually supports the intended puzzle solution until convergence.
Once the layout is verified, the Builder renders a photorealistic image based on the layout and the description. In this final stage, the Examiner again compares the Player’s actions—now based on the image—to the intended solution. If the Player misinterprets object affordances, the Examiner identifies which visual cues are lacking in the image. Using the icon position from the layout stage, we then apply local image editing to enhance or suppress affordances to steer perception toward the correct interaction.
@@ 4 Experiment
We use GPT-4o  [15] API as our base model and generate square images at 1024 $\times$ 1024 resolution with the highest quality. This can be effectively transferred to any blackbox text-to-image model without any specific requirements on model architecture.
@@ 4.1 Evaluation Metrics
We evaluate our method using a combination of human judgments and automated metrics. Specifically, we consider three key human evaluation criteria. Solvability measures whether a player can infer the intended multi-step solution purely from visual cues in the generated image. Shortcut Avoidance assesses whether the image successfully prevents trivial or unintended solutions that bypass the intended logic. Spatial Alignment evaluates how accurately the visual scene reflects the object relations specified in the initial design, such as positions and reachability.
In addition to human assessments, we report the LongCLIP [26] score, which quantifies the semantic alignment between the Designer prompt and the final image, and the average number of image API calls used throughout the generation process (note that this metric is not applicable to baseline methods that lack image-level refinement). Evaluation is conducted on 15 diverse scene settings, each involving 2 core interactive objects. For each case, we collect responses from 10 human annotators, who are shown multiple generated images (from our method and baselines) along with the scene description and object names. They are asked to select the best image for each of the three criteria. We report the percentage of cases where each method’s output is selected as the top choice.
@@ 4.2 Analysis
We compare against vanilla GPT-4o [15] and several ablated versions of our framework by selectively removing refinement stages on the scene graph, layout sketch, and final image. Qualitative results are shown in Fig 3. For example, in the top-right classroom scene with the ladder, naively prompting GPT for an “escape room” (top left) yields a scene with generic visual cues but no coherent solution path. Providing only a textual description introduces some relevant elements but lacks spatial grounding (e.g., a lock floating on the blackboard). Incorporating the scene graph and layout planning significantly improves spatial coherence, aligning the visual arrangement with the intended interaction sequence. Final image refinement further removes visual artifacts (e.g., malformed hooks) and eliminates shortcuts (e.g., climbable desks) that could compromise puzzle integrity.
Quantitative results in Tab 1 reinforce these observations. Starting from vanilla GPT-4o, which achieves only $3.3\%$ solvability and fails entirely on shortcut avoidance, each added module brings measurable gains. The scene graph (+D+S.G.”) more than quadruples shortcut avoidance to $13.3\%$ and introduces modest gains in spatial alignment ($26.7\%$). The addition of layout planning (+D+S.G.+L”) leads to further improvements in both solvability ($10.0\%$) and shortcut avoidance ($20.0\%$), while reducing the average number of required generations. Image-level refinement (“+D+S.G.+I”) helps further suppress unintended interactions and enhances alignment, though at the cost of slightly lower CLIP scores, likely due to localized editing that deviates from the original prompt.
Our full method, combining all stages, achieves the highest scores across all human evaluation metrics: $53.3\%$ for solvability, $46.6\%$ for shortcut avoidance, and $36.7\%$ for spatial alignment—substantially outperforming all baselines. Notably, it achieves these gains with far fewer image generations (4.5 on average), underscoring the efficiency benefits of our structured refinement process. Although CLIP scores slightly decline as editing introduces affordance cues not mentioned in the initial prompt, this trade-off still ensures functional solvability and visual clarity.
@@
5 Human Interface
Figure 4: Real case of human-agent interaction. The human is typing in possible solutions trying to solve the escape room puzzle, while the AI is providing feedback and gradually guiding the player towards the ideal solution. We put them into specific figuring format for better spatial fitting. All texts are direct input and outputs of the VLM conversation.
We develop an interactive human interface that allows users to engage with the generated escape room puzzles by attempting to solve them visually. Alongside the player, an AI agent—equipped with the ground-truth solution—evaluates the human’s proposed actions, providing feedback on whether each step aligns with the intended logic. When discrepancies arise, the agent offers corrective guidance, helping the player iteratively converge toward the correct solution. This setup enables real-time puzzle solving and highlights the interpretability and playability of our generated scenes. See Fig.4 for a walkthrough of the interactive gameplay process.
@@ 6 Conclusion
We presented GenEscape, a hierarchical multi-agent framework for generating 2D escape room puzzles that are both visually realistic and logically solvable. By decomposing the generation process into symbolic reasoning, layout planning, and affordance-aware image refinement, our system ensures functional coherence and visual clarity at each stage. Specialized agents collaborate through iterative feedback loops to simulate human solving and verification, improving puzzle solvability and blocking unintended shortcuts. Experiments across diverse scenes show that our approach significantly outperforms vanilla generation and ablations in both human and automatic evaluations. This work highlights the value of structured, multi-agent coordination for complex visual reasoning tasks.
7 Limitations
@@
Solution to Figure 3
Pool + Lifebuoy
@@
Birthday Party + Balloon
@@
Amusement Park + Seesaw
@@
    Read the sign for the puzzle hint. Observe the message ”BALANCE THE FUN TO ESCAPE!” to infer that equalizing the seesaw is the key to opening the exit.
@@
    Inspect the right side of the seesaw. Notice that it contains immovable 50kg and 20kg weights locked inside a cage, indicating a fixed total weight on that side.
@@
    Collect available weights from the left side. Find a 100kg kettlebell and a golden elephant head placed nearby on the ground.
@@
    Balance the seesaw using the collected items. Carefully place the kettlebell and elephant head on the left side of the seesaw to counterbalance the fixed weights.
@@
    Trigger the hidden mechanism. Once the seesaw achieves balance, pressure on the right end lifts and reveals a trapdoor labeled “EXIT”.
@@
    Escape through the trapdoor. Pull the handle on the EXIT mat to open the trapdoor and leave the room.
Basement + Bomb
@@
    Retrieve the crowbar. Take the crowbar hanging from a wall-mounted shelf.
@@
    Break the glass dome protecting the bomb. Use the crowbar to smash the enclosure and gain access to the bomb.
@@
    Find and collect the matchstick. Look inside an open cabinet nearby and take the matchstick stored inside.
@@
    Ignite the bomb’s fuse. Light the fuse using the matchstick to initiate the detonation sequence.
@@
    Wait for the explosion to create an exit. The bomb detonates and blows a hole through the weakened section of the basement wall.
@@
    Escape through the destroyed wall. Walk through the opening created by the explosion to exit the room.